If a gas in a closed container is pressurized from 15.0 atmospheres to 16.0 atmospheres and its original temperature was 25.0 C, what would the final temperature be?
1
Expert's answer
2020-04-18T07:01:15-0400

n=pV/RT=15 x 1/0.082 x 298=0.614

T2=pV/Rn= 16 x 1 / 0.082 x 0.614 = 318 K. And 45°C